OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -8 - 9t. Find the instantaneous velocity at t = 1 by finding the derivative.

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

find the derivative of s(t)

OpenStudy (anonymous):

ohh ok can you help me with another one please?

OpenStudy (anonymous):

Find the derivative of f(x) = 5x + 9 at x = 2.

OpenStudy (anonymous):

i will try

OpenStudy (anonymous):

5

OpenStudy (anonymous):

derivative if 5x is 5

OpenStudy (anonymous):

Find the derivative of f(x) = 7/ x at x = 1. is this one 7?

OpenStudy (anonymous):

thanks i just wanted to make sure!

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

you welcome